What is abs(3-4i)?
Answer by stanbon(75887)

What is abs(3-4i)
-----
abs(3-4i) = |3-4i| = sqrt(3^2+4^2) = sqrt(25) = 5
================================
Note: |3-4i| means "the distance from (3,0) to (0,-4)"
where 3 is on the x-axis (Real Number) axis and
-4 (Complex Number) is on the y-axis.
================================
Cheers,
Stan H.
==========